Posted before and it was incorrect. Please helpAttempts: 0 Average: 0/1.5 3. Problem 7-03 (Horizon Value of Free Cash Flows) eBook Horizon Value of Free Cash Flows Current
Answer #1

Growth Rate = [FCF(2022) / FCF(2021)] - 1 = [751.75 / 702.567] - 1 = 1.07 - 1 = 0.07, or 7%

Horizon Value = FCF(2022) / [r - g]

= 702.567 / [0.1155 - 0.07] = 702.567 / 0.0455 = 15,442.63, or $15,442.63 million